Grade 11Discuss with colleagues and IITians

A point p(x,y) moves so that the sum of the distances from P to the coordinate axes is equal to the distance from P to the point A(1,1). The equation of the locus of P in the first quadrant is
(a) (x+1)(y+1)=1
(b) (x+1)(y+1)=2
(C) (x-1)(y-1)=1
(D) (x-1)(y-1)=2

Raviraj singh
11 Years agoGrade 11

1 Answer

rajesh
11 Years ago
let P be (x,y)
as per the question sum of the distances of P from the coordinate axes is 
x+y
(x+y)2=(x-1)2+(y-1)2
x2+y2+2xy=x2+1+y2+1-2x-2y
2x+2y+2xy=2
x+y+xy=1
x+y+xy+1=2
x+1+y+xy=2
(x+1)+y(x+1)=2
(x+1)(y+1)=2